OpenStudy (anonymous):

7/17 = x/100 10/17 = x/100 how do I solve ?

OpenStudy (anonymous):

how do i solve for x?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let's look at 7/17 first

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

ok, can you do long division?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

one way to isolate x is multiply both sides by 100 --> (7/17) * 100 = x / 100 * 100. therefore the left side is (7/17)*100 = x

OpenStudy (anonymous):

so you do cross multiplication? wait. we have to do persentages of our class 7 girls and 10 guys for A total of 17 so how do i do that?

OpenStudy (anonymous):

ok, so you can solve for 7/17. that would yield a decimal once you do the division. once you have the decimal, you then multiply that value by 100 to obtain the percentage

OpenStudy (anonymous):

7/17 = .411767....

OpenStudy (anonymous):

okay i got the .411764706

OpenStudy (anonymous):

you can see that 7/17 + 10/17 = 17/17 which is 1. which would be 100%. Here we look at 7/17 which is .411767, which is 41.1767%. you see that multiplying .411767 by 100 moves the decimal point to the right

OpenStudy (anonymous):

yes is do understand that.

OpenStudy (anonymous):

do you think you can to solve for 10/17 as a percentage?

OpenStudy (anonymous):

yes thank you so much!

OpenStudy (anonymous):

cool, no problem